A project with a 3 year life and a cost of R95 000 generates revenues of R32 000 in year 1, R44 000 in year 2, and R52 000 in year 3. If the discount rate is 14%, what is the NPV of the project?

1
Expert's answer
2022-01-16T13:31:49-0500

Project life = 3 years

Initial investment = R95,000

Revenues in year 1 = R32,000

Revenues in year 2 = R44,000

Revenues in year 3 = R52,000

Discount rate = 14%

The NPV of the project is:


"NPV=\Sigma\frac{CF_n}{(1+i)^n}-I_0"


"NPV=\frac{32000}{(1+0.14)^1}+\frac{44000}{(1+0.14)^2}+\frac{52000}{(1+0.14)^3}-95000"


"NPV=\frac{32000}{(1.14)^1}+\frac{44000}{(1.14)^2}+\frac{52000}{(1.14)^3}-95000"


"NPV=28064+33836+35100-95000"


NPV = R2,000